Let f be one of the following functions on R. Compute f-hat. (a) f(x)=cosx, xER. (b) f(x)=sinx, xER. (c) f(x)=e^(iax), xER, where aER.
1
Expert's answer
2019-11-12T13:00:53-0500

Fourier transform of function is:

"\hat{f}(k) = \frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty} dx\, e^{ikx}f(x)"

and integral representation of Dirac delta-function is

"\delta(x)=\frac{1}{2\pi}\int_{-\infty}^{\infty}e^{ikx}\, dk"

a)

"f(x)= \cos(x)\\
\hat{f}(k) = \frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty} dx\, e^{ikx}\cos(x) = \frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dx\, e^{ikx}\frac{e^{ix}+e^{-ix}}{2}=\\
\sqrt{\frac{\pi}{2}}\delta(k-1)+\sqrt{\frac{\pi}{2}}\delta(k+1)"

b)

"f(x)= \sin(x)\\
\hat{f}(k) = \frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty} dx\, e^{ikx}\sin(x) = \frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dx\, e^{ikx}\frac{e^{ix}-e^{-ix}}{2i}=\\
i\sqrt{\frac{\pi}{2}}\delta(k-1)-i\sqrt{\frac{\pi}{2}}\delta(k+1)"

c)

"f(x)= e^{iax}\\
\hat{f}(k) = \frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty} dx\, e^{ikx}e^{iax} =\\
 =\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty} dx\, e^{i(k+a)x} = \sqrt{2\pi}\delta(k+a)"